OpenStudy (anonymous):

Name the binomial you can multiply by (x + 9) to get the product x2 + 5x - 36.

OpenStudy (anonymous):

x2 - 5x - 36 = 0 (x - 9) (x + 4) = 0 x = 9 and x = -4 does this seem right?

OpenStudy (anonymous):

yes, that is correct

OpenStudy (anonymous):

Thanks! :)